Reports 1-1 of 1 Clear search Modify search
mark.barton - 11:07 Tuesday 22 November 2016 (2176) Print this report
BS Test Hang

Monday, with Enzo.


  • We did a bunch of fixed-sine actuation tests from input at the COILOUTF blocks to output of the OSEMINF blocks, to check the polarity of the actuation and find what gain values (+1 or -1) to put in the COILOUTF blocks. The frequency was chosen as 0.3 Hz so as to be below all the suspension resonances, and all the COILOUTF gains (except H1, by accident) were set to +1 to begin with. In fact for the phases we got:
    • V1: 0°
    • V2: -179°
    • V3: -179°
    • H1: 11° (had -1 in COILOUTF)
    • H2: -178°
    • H3: -175°
  • The data is in /kagra/Dropbox/Subsystems/VIS/TypeBData/BSTestHang/IRM/2016-11-21-BSTH-IRM-DTT-IM-??-01.xml.
  • All the magnets had N facing out except V1, so it looks like all channels except V1 will need -1's in COILOUTF.
  • We fixed various problems with MEDM screens:
    • We created a new IM watchdog screen VIS_CUST_TYPEB_IM_WD_OSEMAC_BANDLIM.adl (the existing one had been a copy of the TM version).
    • We copied band limit filters to IM from TM.
    • We noted but did not fix a problem with the TM version VIS_CUST_TYPEB_TM_WD_OSEMAC_BANDLIM.adl, which should be getting the raw OSEM signals as input but is getting something else.
    • We fixed the VIS_CUST_TYPEB_IM_CENTERING_??ZOOM.adl screens which were not getting called correctly by VIS_CUST_TYPEB_IM_CENTERING.adl (the macro file arguments had "sus" instead of "vis" in several places).
    • We corrected a bunch of screens that had LIGO-style channel names (F1, F2, F3 T1, T2, SD and the like) and other cosmetic errors.
Search Help